INFRARED SPECTRA OF C_{2}H_{4} UNDER JET-COOLED AND PARA-H_{2} MATRIX CONDITIONS

\underline{YUAN-PERN LEE} AND CHUNG-WEI HUANG, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan and Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan; KEEVIN DIDRICHE^{a} AND MICHEL HERMAN, Laboratoire de Chimie quantique et Photophysique, CP160/09, Université libre de Bruxelles, Ave. Roosevelt, 50, B-1050, Brussels, Belgium; MAUD ROTGER, Institut Carnot de Bourgogne, UMR 5209 CNRS-Université de Bourgogne, 9 avenue Alain Savary, B.P. 47870, F-21078 Dijon Cedex, France.

Spectra of C_{2}H_{4} isolated in para-H_{2} are compared with those of jet-cooled C_{2}H_{4} recorded with an FTIR spectrometer. The ν_{7}, ν_{9}, and ν_{11} bands of jet-cooled C_{2}H_{4} all exhibit an intensity distribution corresponding to about 10 K for rotation, with no evidence of nuclear spin conversion. Spectra of the ν_{11} band of C_{2}H_{4} isolated in solid para-H_{2} clearly show a rotational structure similar to that observed for C_{2}H_{4} in helium droplets. Other bands of C_{2}H_{4} exhibit only partially resolved rotational structures. Rotational constants B and C for ν_{11} of C_{2}H_{4} in para-H_{2} are about 50 % those of the gaseous C_{2}H_{4}. The out-of-plane deformation mode ν_{7} shows a small (< 1 cm^{−1}) blue matrix shift, whereas other modes are red-shifted by 2.2 to 3.9 cm^{−1}.

aFRIA Researcher